OpenStudy (anonymous):

Help Me Please : ) A number cube with faces labeled 1–6 is rolled 72 times. About how many times would the result be 4? A. 4 B.6 C.8 D.12

OpenStudy (abmon98):

72/6=12

OpenStudy (anonymous):

A bag contains a total of 25 marbles. Five of the marbles are blue, 3 are red, 8 are green, and 9 are white. If a marble is drawn from the bag and then replaced 50 times, about how many times would a green marble be drawn? A.4 B.8 C.16 D.24

OpenStudy (abmon98):

3+8+5+9=25, 8/25*50=16

OpenStudy (anonymous):

Scott plays a game using the spinner shown. During the game he will spin a total of 42 times. Which is the best prediction for the number of times he will get either "back 2" or "back 3" on the spinner? A.7 B.14 C.21 D.28 http://static.k12.com/calms_media/media/1416500_1417000/1416596/1/709abf9954e37e8a3d76e2656348ce89eecf1619/FGA_130913_160912.jpg thats the URL